OpenStudy (anonymous):

how do i do this?? 21% of 50?

OpenStudy (anonymous):

0.21% x 50 = 10.5%

OpenStudy (anonymous):

u put a . infront of of the 21? do u do this on every problem thats the same thing?

OpenStudy (anonymous):

Yes, if you are doing percentages. But, remember if it's a number less than 10 put a 0 after the decimal point and before the number. EX: 5% would be 0.05% NOT 0.50%!!!

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

so 45%of 72 would be 32.4?

OpenStudy (anonymous):

Yeah. 0.45% x 72 = 32.4%

OpenStudy (anonymous):

would 140% of 85 be 119?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

ok i kno how to do that now but ive got a diff type of problem to deal with... 86 is 12% of what number??

OpenStudy (anonymous):

86 / x = 12% 86 / x = 0.12 86 / 0.12 = x x = 716.666667 or 716.7

OpenStudy (anonymous):

alright thank you lots for your time and help

OpenStudy (anonymous):

NP :)